# A way around the exponential scaling in optimal quantum control

### v1

@@ II.3.3 Numerical optimization
In order to be evaluable without exponential effort, it is essential that the state vector |$\ket{\psi(T)}$ does not need to be constructed explicitly; but since the expectation values $\bra{\psi(T)}Q\ket{\psi(T)}$ can be constructed with the initial state |$\ket{\Psi(0)}$ and a time-evolved operator $Q(T)$, no such state construction is necessary. The propagation of $Q(t)$ follows Eq. (8), and is thus not impeded by the exponential growth of the Hilbert space. As long as expectation values with respect to the initial state |$\ket{\Psi(0)}$ can be specified efficiently, as it is typically the case in problems of state preparation, adiabatic quantum computation or quantum approximate optimization algorithms [37, 38, 39], the state fidelity $F$ can be bounded efficiently by Eq. (19).
III Diabatic state transfer in a driven spin chain
@@
The approach outlined in Sec. II above is applicable to any set of operators satisfying Eq. (5). The benefits of this approach depend on the size $d$ of the resulting set $\{\mathfrak{a}_{j}\}$. Sec. V provides a list of exemplary spin systems and sets of operators with a size that is polynomial in the number of spins. Since such a list of examples is necessarily a bit technical, the following discussion first illustrates solutions to explicit control problems following the general framework outlined in Sec. II with a linear spin chain.
The utilised system Hamiltonian [40]
$H(t)=\sum_{j=1}^{n}f_{j}(t)Z_{j}+g\sum_{j=1}^{n-1}X_{j}X_{j+1}+\sum_{j\in\{1,n\}}h_{j}(t)X_{j}$ (20)
contains nearest-neighbor interactions in terms of the Pauli $X$ operators, and single-spin energies in terms of the Pauli $Z$ operators and the Pauli $X$ operators on the end-spins of the chain (see Fig. 1).
@@
The Hamiltonian $H(t)$ in Eq. (20), on the other hand, has no conserved quantities for $g\neq 0$ and general time-dependent functions $f_{j}(t)$ and $h_{j}(t)$; i.e., there is no operator $A$ (apart from multiples of the identity) that commutes with all operators $Z_{j}$, $X_{1}$, $X_{n}$ and $\sum_{j=1}^{n-1}X_{j}X_{j+1}$ as discussed in more detail in Appendix A.
@@
The present framework thus reduces a control problem in an exponentially large Hilbert space to a problem in a vector space of quadratic growth. Thanks to the quadratic scaling, it is possible to design optimal control protocols for spin chains of lengths that are inaccessible for approaches based on quantum states. The following examples are motivated by typical problems of quantum simulation and state preparation or stabilisation, and they exploit the fact that the Lie algebra given in Eq. (21j) contains three-body interactions such as $X_{j}Z_{j+1}X_{j+2}$ or even $n$-body interactions that cannot be realised by static means.
@@ III.1 Target Hamiltonians
Following the discussion of Sec. II.3.1, one can seek to find time-dependent functions $f_{j}(t)$ and $h_{j}(t)$ in the system Hamiltonian $H(t)$ (Eq. (20)), such that the system initialized in the ground-state of a non-interacting Hamiltonian evolves towards the ground-state of a Hamiltonian with interactions.
A natural choice of non-interacting Hamiltonian based on Eq. (20) is given by the initial conditions $f_{j}(0)=f_{k}(0)$ for all $j,k$, $f_{j}(0)\gg g$ and $h(0)=0$, i.e. non-interacting spins with single-spin-Z Hamiltonian.
@@
whose ground state is a GHZ state, which holds significant importance in the fields of quantum sensing [42], quantum communication [43, 44], and macroscopic quantum mechanics [45], attracting great experimental interest [46, 47]. Another interesting example is the Hamiltonian
@@
whose ground state is the one-dimensional $n$-qubit cluster state for measurement-based quantum computation [48]. Since the Hamiltonians $H_{G}$ and $H_{C}$ admit an analytic construction of a target invariant, the following discussion includes also the example
@@ III.2 Control targets
As discussed in Sec. II.3.1, a natural choice for the initial condition ${\cal I}(0)$ is given by
@@
The definition of corresponding control target, however requires some care, and it needs to be done separately for the three different control problems defined in terms of the three Hamiltonians $H_{G},H_{C}$ and $H_{D}$.
@@ III.2.3 Control target for $H_{D}$
While control targets could be defined analytically for the control problem of realizing the ground state of $H_{G}$ and $H_{C}$, the case of $H_{D}$ seems to require a numerical construction.
@@
$H_{a}(t)=\left(1-\frac{t}{T_{a}}\right)H_{0}+\frac{t}{T_{a}}H_{D}+\frac{t}{T_{a}}\left(1-\frac{t}{T_{a}}\right)H_{B}\ ,$ (30)
@@
vanishes at the initial and final point in time ($t=0$ and $t=T_{a}$), but turns some level crossings into avoided crossings, so that adiabatic dynamics does indeed ensure ground-state-to-ground-state transfer.
A control target ${\cal I}_{T}={\cal I}_{a}(T_{a})$ can be constructed numerically with the adiabaticity parameter $T_{a}$ sufficiently large. Since any numerical propagation is based on Eq. (8), the numerical effort follows the quadratic scaling of the Lie algebra (Eq. (21j)) and not the exponential scaling of the Hilbert space.
@@ III.3 Results – State transfer
Figure 2: Infidelity for dynamics for a spin chain (Eq. (20)) with optimized single-qubit driving, as function of systems size $n$. Solid shapes depict infidelities $\cal J$ (Eq. (12)) in log-scale for state transfer (panel a) and infidelities ${\cal J}_{U_{T}}$ (Eq. (14)) for quantum gates (panel b). Empty diamonds and squares in panel a depict bounds (Eq. (19)) on state infidelities; empty triangles in panel a represent state infidelities and empty shapes in panel b depict gate infidelities. As result of the optimization, the values of the infidelities $\cal J$ and ${\cal J}_{U_{T}}$ are largely independent of the system size $n$. Actual state- and gate-infidelities do grow with increasing system size, but they remain well below what can be achieved given the experimental imperfections of actual quantum devices.
@@
Fig. 2a depicts the infidelity ${\cal J}$ (Eq. (12)) obtained with numerically optimized, time-dependent Hamiltonians as function of the number $n$ of spins in a chain with $n$ ranging from $5$ to $50$. The different solid symbols correspond to the three different target operators discussed in Sec. III.1, namely $H_{G}$ (Eq. (22)) (red diamonds), $H_{C}$ (Eq. (23)) (blue squares) and $H_{D}$ (Eq. (24)) (black triangles). The optimisations for $H_{G}$ and $H_{C}$ are intentionally terminated when infidelities below $10^{-6}$ are reached. The operator $H_{D}$, however has a vanishing gap in the limit $n\to\infty$, and the closing gap requires longer running time; in the case of $H_{D}$, the optimisation is thus terminated once a maximum number of iterations, around $10^{3}$, is reached.
The empty red diamonds and blue squares depict the bound of Eq. (19) on the state fidelity for the ground states of $H_{G}$ and $H_{C}$. The closing gap of $H_{D}$ on the other hand makes this bound unsuited for $H_{D}$. The state fidelity in this case is thus estimated from numerical propagations with matrix product states [49, 50] for qubit numbers $n\leq 30$, as depicted by the empty triangles in Fig. 2a. Despite the efficient matrix-product representation of the initial and final states, the states at in-between times created by the controlled dynamics for $n\geq 35$ require bond dimensions exceeding $500$, rendering numerical optimisation impractical.
@@
In all three cases shown in cases Fig. 2a, the achieved state fidelities confirm successful optimal control for system sizes far outside the regime accessible to explicit state control. In particular, the state infidelities are all well below the limits imposed by system imperfections of current or foreseeable quantum devices [51].
@@
Interestingly, because of the constant interaction term in the system Hamiltonian (Eq. (20)) there is a minimal time required for the state transfer, also referred to as quantum speed limit [52]. In order to obtain infidelities as low as in Fig. 2, the duration of the controlled dynamics needs to increase with the system size, but a modest approximately linear increase ($T\sim n\pi/2g$) with the qubit number $n$, and the timescale $1/g$ defined by the interaction constant $g$ in Eq.(20) is sufficient for the ground state transfer.

### v2

@@ II.4.2 State fidelity
The state infidelity $1-F$ can thus be bounded in terms of the ground state energy $E_{0}$, the first excited state energy $E_{1}$ of the Hamiltonian $H_{T}$ and the expectation value $\bra{\Psi(T)}H_{T}\ket{\Psi(T)}$. At first sight, it might seem that evaluating this expectation value requires explicit construction of the state vector |$\ket{\Psi(T)}$. Since, however, the state vector |$\ket{\Psi(t)}=U(t)\ket{\Psi(0)}$ follows the unitary dynamics $U(t)$ induced by the system Hamiltonian, the expectation value can be evaluated in terms of the initial state |$\ket{\Psi(0)}$, and a time-evolved operator $U^{\dagger}(T)H_{T}U(T)$. The latter can be obtained by backward-propagation of Eq. (1), and for any operator $H_{T}$ expanded in terms of the operators $\{\mathfrak{a}_{j}$} ($j\in[1,d]$) this can be done with same effort as required for propagation of ${\cal I}(t)$. As long as expectation values with respect to the initial state |$\ket{\Psi(0)}$ can be specified efficiently, as it is typically the case in problems of state preparation, adiabatic quantum computation or quantum approximate optimization algorithms [39, 27, 40], the state fidelity $F$ can be bounded efficiently by Eq. (18).
II.4.3 Desired unitary
Similar to the problem of state preparation, the task of realizing a specific unitary – the central problem in quantum simulation [41] – can also be addressed using the present approach, though there are some key differences. In the above problem of state preparation, there is not a unique choice of control target ${\cal I}_{T}$ for a given initial condition ${\cal I}(0)$. This is different for a control problem with the goal to realize a given unitary ${\cal U}_{T}$. In this case, the control target corresponding to an initial condition ${\cal I}(0)$ necessarily needs to read ${\cal I}_{T}={\cal U}_{T}{\cal I}(0){\cal U}_{T}^{\dagger}$.
If there is a unitary $V$ that commutes with ${\cal I}_{T}$, i.e. $[V,{\cal I}_{T}]=0$, then also the relation ${\cal I}_{T}=V{\cal I}_{T}V^{\dagger}=V{\cal U}_{T}{\cal I}(0){\cal U}_{T}^{\dagger}V^{\dagger}$ holds. Reaching the target ${\cal I}_{T}$ does thus not necessarily verify that the propagator ${\cal U}_{T}$ is realized. Ensuring that one operator ${\cal I}(t)$ with an initial condition ${\cal I}(0)$ evolves towards the target ${\cal I}_{T}={\cal U}_{T}{\cal I}(0){\cal U}_{T}^{\dagger}$ that is consistent with the unitary ${\cal U}_{T}$, therefore, does not necessarily guarantee that the unitary ${\cal U}_{T}$ is indeed realized. Similarly to the definition of a faithful objective function in terms of a set of initial and final states, also the present approach requires a set of initial conditions ${\cal I}_{j}(0)$ and corresponding targets ${\cal I}_{Tj}={\cal U}_{T}{\cal I}_{j}(0){\cal U}_{T}^{\dagger}$ for the definition of an objective function that is minimized exactly for the desired unitary ${\cal U}_{T}$. Unlike standard methods for gate characterization, which often require an exponentially large set of initial and final states [42, 43], this operator-based approach typically requires only a small number of initial conditions [44].
A possible choice of a faithful objective function is given by
with a sufficiently large number $N_{c}$ of initial conditions $\mathcal{I}_{j}(0)$. The initial conditions $\mathcal{I}_{j}(0)$ can, for example, always be chosen as a minimal set of operators from $\{\mathfrak{a}_{j}\}$ such that there is no operator in $SU(2^{n})$ that commutes with all the transformed operators $\mathcal{U}_{T}\mathcal{I}_{j}(0)\mathcal{U}_{T}^{\dagger}$.
Since each summand in Eq. (19) can be evaluated analogously to Eq. (12), the objective function ${\cal J}_{{\cal U}_{T}}$ can indeed be evaluated with polynomial effort.
II.4.4 Numerical optimization
Once a control target ${\cal I}_{T}$ and an objective function, such as ${\cal J}$ (Eq. (12)) or ${\cal J}_{{\cal U}_{T}}$ (Eq. (19)) is specified, one can use any numerical pulse shaping algorithm [1] to design a time-dependent Hamiltonian that realizes a dynamics from ${\cal I}(0)$ to ${\cal I}(T)$, at a final time $T$, that is close to ${\cal I}_{T}$. Since Eq. (8) with an anti-symmetric matrix $K$ induces orthogonal (unitary and real) dynamics, any numerical algorithm that is applicable to the Schrödinger equation (such as Krotov [45], GRAPE [13] or CRAB [16]) is directly applicable, and techniques to construct e.g. gradients with respect to control parameters also directly apply to Eq. (8) (see Appendix B for more details on the implementation).
II.5 A birdseye view on the algorithm
Putting everything together, the resultant methodology for state-control can be summarized in terms of the following algorithm.
Given a Hamiltonian with tuneable time-dependence (Eq. (3)) and a set of operators $\{\mathfrak{a}_{j}\}$, such that the closed commutator relationships (Eq. (5)) are satisfied,
    choose an initial condition ${\cal I}(0)=\sum_{j}a_{j}(0)\mathfrak{a}_{j}$, with a non-degenerate ground state.
    find a control target ${\cal I}_{T}$ that implicitly specifies the target state, and that needs to be reachable from initial condition ${\cal I}(0)$ with the dynamics induced by the system Hamiltonian.
    design an optimized time-dependence of the system Hamiltonian (Eq. (3)), with the dynamics described by Eq. (8), to minimize an objective function formulated in terms of the expansion coefficients of the control target ${\cal I}_{T}$ in the set of operators $\{\mathfrak{a}_{j}\}$. The design of this time-dependence can be pursued with any pulse-shaping algorithm that applies to linear differential equations.
The resultant time-dependent Hamiltonian is such that the system, if initialized in the ground state of ${\cal I}(0)$, evolves towards the ground state of ${\cal I}_{T}$.
The algorithm for control of unitary dynamics is largely equivalent, but it requires more that one initial condition, and the target for each initial condition is defined uniquely through the desired unitary.
III State preparation in a driven spin chain
@@
The approach outlined in Sec. II above is applicable to any set of operators satisfying Eq. (5). The benefits of this approach depend on the size $d$ of the resulting set $\{\mathfrak{a}_{j}\}$. There does not seem to be an existing algebraic framework to determine the system Hamiltonians with an associated set of operators $\{\mathfrak{a}_{j}\}$ that is significantly smaller than the size of the system Hilbert space, but the existence of such Hamiltonians can be demonstrated with explicit examples.
Sec. V provides a list of exemplary spin systems and sets of operators with a size that is polynomial in the number of spins. Since such a list of examples is necessarily a bit technical, the following discussion and the subsequent section Sec. IV will illustrate the application of the general control framework outlined in Section II with the specific case of a linear spin chain with nearest-neighbor interactions, as depicted in Figure 1.
In such a configuration, the system dynamics is induced by the system Hamiltonian
$H(t)=g\sum_{j=1}^{n-1}X_{j}X_{j+1}+\sum_{j=1}^{n}f_{j}(t)Z_{j}+\sum_{j\in\{1,n\}}w_{j}(t)X_{j}\ ,$ (20)
which consists of a constant interaction term and single-qubit control terms. The constant term contains nearest-neighbor interactions in terms of the Pauli $X$ operators, while the single-qubit control terms, with tuneable time-dependent functions $f_{j}(t)$ and $w_{j}(t)$, contain the Pauli $Z$ operators and the Pauli $X$ operators on the end-spins of the chain (see Fig. 1).
@@
The Hamiltonian $H(t)$ in Eq. (20), on the other hand, has no conserved quantities for $g\neq 0$ and general time-dependent functions $f_{j}(t)$ and $w_{j}(t)$; i.e., there is no operator $A$ (apart from multiples of the identity) that commutes with all operators $Z_{j}$, $X_{1}$, $X_{n}$ and $\sum_{j=1}^{n-1}X_{j}X_{j+1}$ as discussed in more detail in Appendix A.
@@
While the system Hamiltonian (Eq. (20)) contains only the terms $Z_{j}$, $X_{j}X_{j+1}$, $X_{1}$ and $X_{n}$, the operator ${\cal I}(t)$ needs to be expanded in terms of the full set of Eq. (21j) so that Eq. (5) is satisfied and the Ansatz in Eq. (4) is sufficiently general to allow for integration of Eq. (1).
Since Eq. (21j) contains a quadratic number of terms, the present framework reduces a control problem in an exponentially large Hilbert space to a problem in a vector space of quadratic growth. Thanks to this quadratic scaling, it is possible to design optimal control protocols for spin chains of lengths that are inaccessible for approaches based on quantum states.
The following examples are motivated by typical problems of quantum simulation and state preparation or stabilisation. They exploit the fact that the Lie algebra given in Eq. (21j) contains three-body interactions such as $X_{j}Z_{j+1}X_{j+2}$ or even $n$-body interactions that cannot be realised by static means. The goal of the following two sections Sec. III.1 and IV.1 will thus be to find time dependent functions $f_{j}(t)$ and $w_{j}(t)$ in the system Hamiltonian $H(t)$ (Eq. (20)) that ensure that the system dynamics with only the two-body interaction in Eq. (20) is helpful for preparation of states or realization of propagators that are characterized by three-body interactions or interactions between more than three spins.
@@ III.1 Target Hamiltonians
Following the discussion of Sec. II.4.1, one can seek to find time-dependent functions $f_{j}(t)$ and $w_{j}(t)$ in the system Hamiltonian $H(t)$ (Eq. (20)), such that the system evolves from an initial state to a desired final state. Typical choices for an initial state are given by separable states that can be specified efficiently, but final states of interest are often highly entangled states. The following discussion will thus avoid the explicit specification of a state vector, but both the initial state and the final state to be realized will be specified in terms of a Hamiltonian to which they are the ground state.
A natural choice of non-interacting Hamiltonian based on Eq. (20) is given by the initial conditions $f_{j}(0)=f_{k}(0)$ for all $j,k$, $f_{j}(0)\gg g$ and $w_{j}(0)=0$ for all $j$, i.e. non-interacting spins with single-spin-Z Hamiltonian.
Crucially, the Hamiltonians defining the final state to be realized do not need to be of the form of the system Hamiltonian in Eq. (20), but any Hamiltonian comprised of the operators given in Eqs. (21j) including three-spin interactions or interactions between more than three spins is a viable choice.
@@
whose ground state is a GHZ state, which holds significant importance in the fields of quantum sensing [47], quantum communication [48, 49], and macroscopic quantum mechanics [50], attracting great experimental interest [51, 52].
Another interesting example is the Hamiltonian
@@
whose ground state is the one-dimensional $n$-qubit cluster state for measurement-based quantum computation [53].
Since the Hamiltonians $H_{G}$ and $H_{C}$ admit an analytic construction of a target invariant, the following discussion includes also the example
@@ III.2 Control targets
As discussed in Sec. II.4.1, a natural choice for the initial condition ${\cal I}(0)$ is given by
@@
The definition of corresponding control target, however, requires some care, and it needs to be done separately for the three different control problems defined in terms of the three Hamiltonians $H_{G},H_{C}$ and $H_{D}$.
Figure 2: Infidelity for dynamics for a spin chain (Eq. (20)) with optimized single-qubit driving, as function of systems size $n$. Solid shapes depict operator infidelities $\cal J$ (Eq. (12)) in log-scale for state transfer (panel a) and operator infidelities ${\cal J}_{U_{T}}$ (Eq. (19)) for quantum gates (panel b). Empty diamonds and squares in panel a depict bounds (Eq. (18)) on state infidelities; empty triangles in panel a represent state infidelities based on matrix-product-state representations and empty shapes in panel b depict gate infidelities. As result of the optimization, the values of the infidelities $\cal J$ and ${\cal J}_{U_{T}}$ are largely independent of the system size $n$. Actual state- and gate-infidelities do grow with increasing system size, but they remain well below what can be achieved given the experimental imperfections of actual quantum devices.
@@ III.2.3 Control target for $H_{D}$
While control targets could be defined analytically for the control problem of realizing the ground state of $H_{G}$ and $H_{C}$, the case of $H_{D}$ seems to require a numerical construction. The target $\mathcal{I}_{D}=H_{D}$ is unreachable from $\mathcal{I}(0)$ with unitary dynamics, so a numerical target $\mathcal{I}_{D}$ with the same ground-state as $H_{D}$ must be constructed.
@@
$H_{a}(t)=\left(1-\frac{t}{T_{a}}\right)H_{I}+\frac{t}{T_{a}}H_{D}+\frac{t}{T_{a}}\left(1-\frac{t}{T_{a}}\right)H_{B}\ ,$ (30)
@@
vanishes at the initial and final point in time ($t=0$ and $t=T_{a}$), but turns some level crossings into avoided crossings, so that adiabatic dynamics does indeed ensure transfer from the ground-state of $H_{I}$ to the ground-state of $H_{D}$.
A control target ${\cal I}_{D}={\cal I}_{a}(T_{a})$ can be constructed numerically with the adiabaticity parameter $T_{a}$ sufficiently large. Since any numerical propagation is based on Eq. (8), the numerical effort follows the quadratic scaling of the Lie algebra (Eq. (21j)) and not the exponential scaling of the Hilbert space.
@@ III.3 Results – State transfer
Fig. 2a depicts the infidelity ${\cal J}$ (Eq. (12)) obtained with numerically optimized, time-dependent Hamiltonians as function of the number $n$ of spins in a chain with $n$ ranging from $5$ to $50$. The different solid symbols correspond to the three different target operators discussed in Sec. III.1, namely $H_{G}$ (Eq. (22)) (red diamonds), $H_{C}$ (Eq. (23)) (blue squares) and $H_{D}$ (Eq. (24)) (black triangles). The optimisations for $H_{G}$ and $H_{C}$ are intentionally stopped when infidelities fall below $10^{-6}$. The optimisation for $H_{D}$, however, requires longer running time and is therefore terminated once a maximum number of iterations, around $10^{3}$, is reached.
The empty red diamonds and blue squares depict the bound of Eq. (18) on the state fidelity for the ground states of $H_{G}$ and $H_{C}$. The gap between the infidelities for the operators and the bounds on the state infidelity grows approximately linearly with the number of qubits (for more details, see Appendix C). For practical implementations, this growing gap does not seem to pose a relevant limitation, since state infidelities on the order of $10^{-5}$ are below what is currently achievable or verifiable even from few-qubit states [54].
The Hamiltonian $H_{D}$ is gap-less, i.e., the gap between the ground state and the first excited state vanishes in the thermodynamic limit. For finite-size chains, this gap gets progressively smaller with increasing system size, making the bound of Eq. (18) unsuited for $H_{D}$. The state fidelity in this case is thus estimated from numerical propagations with matrix product states [55, 56] for qubit numbers $n\leq 30$, as depicted by the empty triangles in Fig. 2a. Despite the efficient matrix-product representation of the initial and final states, the states at in-between times created by the controlled dynamics for $n\geq 35$ require bond dimensions exceeding $500$, rendering numerical propagation impractical. Both operator infidelities (Eq. (12)) and state infidelities are larger than for $H_{G}$ and $H_{C}$. This is mostly due to the closing gap of $H_{D}$ that results in increasing errors in the adiabatic evolution used to obtain the target $\mathcal{I}_{D}$. The optimization of the control pulses is made consistent with the estimated error of $\mathcal{I}_{D}$, i.e. the optimization algorithm is terminated when infidelities of the order of the accuracy of $\mathcal{I}_{D}$ are reached. For any finite system, this error can be reduced further if needed, and the present accuracy is chosen in accordance with the practical limitations of existing quantum devices [54].
Figure 3: Optimised pulses, $f_{1}$, $f_{25}$ and $f_{50}$, for state transfer in a chain with 50 spins as function of time $t$ in multiples of the interaction time $\tau_{g}=2\pi/g$. The operator infidelities obtained with these pulses are $\approx 10^{-5}$ for the GHZ state (a), $\approx 10^{-5}$ for the cluster state (b), and $\approx 2\times 10^{-5}$ for the ground state of $H_{D}$ (c). The duration of controlled dynamics is longer for $H_{D}$ in (c) than in the other cases. This is likely due to the difficulty in realising the ground state of a gap-less model.
@@
In all three cases shown in cases Fig. 2a, the achieved state fidelities confirm successful optimal control for system sizes far outside the regime accessible to explicit state control. All deviations from perfect, vanishing infidelity are due to finite numerical accuracy. In Fig. 2, the accuracy is chosen to not exceed the limits imposed by system imperfections of current or foreseeable quantum devices too excessively, but higher accuracies can be reached if technological advances require more accurate solutions.
@@
Interestingly, because of the constant interaction term in the system Hamiltonian (Eq. (20)), there is a minimal time required for the state transfer, also referred to as quantum speed limit [57]. In order to obtain infidelities as low as in Fig. 2, the duration of the controlled dynamics needs to increase with the system size, but a modest approximately linear increase $\sim n\tau_{g}/4$ for $H_{G}$ and $H_{C}$, and $\sim n\tau_{g}/2$ for $H_{D}$, with the qubit number $n$ and the timescale $\tau_{g}=2\pi/g$ given by the interaction constant $g$ in Eq. (20), is sufficient for the ground state transfer. The time-dependence in the system Hamiltonian is parametrized in terms of piecewise constant pulses that are suitable for numerical optimization algorithms. To ensure the linear increase in gate-time while maintaining a constant control duration per time bin across all system sizes, the number of time bins in the pulse must also increase linearly with the system size. The results in Fig. 2 are obtained with pulses of $10n$ bins, except for $U_{G}$ with $20n$ bins, but the exact number can be adjusted depending on the desired infidelity.
Despite the challenging goal of realizing eigenstates of Hamiltonians beyond pairwise interactions, the necessary control pulses do not exhibit undesired characteristics like excessively high amplitudes or frequencies. Even optimizations without penalty for high amplitudes do not yield amplitudes exceeding the coupling strength $g$ by a factor of $2$. The spectral width obtained with a standard optimization (see Appendix B.3) is not excessive, but it is broader than necessary. It is thus possible to smooth pulses with only a moderate increase in infidelity, and to use these smoothed pulses as initial condition for further optimization.
Fig. 3 depicts pulse shapes obtained in such a fashion for $f_{1}(t)$, $f_{25}(t)$ and $f_{50}(t)$ in a chain with $n=50$ spins, i.e. driving patterns for the spins at the edges and center of the chain. In addition to the smooth temporal shape, the highest amplitudes are of the order of $g$, and even this moderately strong driving arises only in short intervals.
For the GHZ state, the optimised pulses hint at a pattern of sequential driving of the spins from left to right along the chain, as the pulses for the spins on the left peak first. This is shown in Fig. 3a for the 1st, 25th and 50th spins, and this sequential pattern does indeed occur for the pulses of the other spins in the chain.
The pulse shapes for smaller chains have qualitatively similar shapes, but both the duration and the number of time bins are approximately proportional to the system size. The pulse data for all spins and system sizes ranging from 5 to 50 is provided in Ref. [58].